40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Engineering group Lab elog, Page 5 of 6  Not logged in ELOG logo
ID Date Author Type Category Subject
  63   Tue May 22 15:08:04 2018 Calum, Luis MechanicsElectric Field MeterEfm he test unit 002 test take 2

Initial pump down was 3.5e-10 torr l/s with a pressure 7.1 e-3 torr.


after 1 minute it reachs 5.6 e-10 torr l/s  then continues to change until the pressure hits (at 2 minutes)  6.2 e-10 torr l/s ;

the pressure stay steady at the 2:40minutes mark then gradually start to change, 6.0e-10torr l/s was seen at the 3 minutes mark.

after the 3 minutes, we read 6.0e-10 torr l/s with a pressure of  6.9e-3 torr.

 

 

Attachment 1: EFM2_Helium_Leak_Test.mov
Attachment 2: EFM2_Helium_Leak_Test.mp4
  62   Tue May 22 14:13:02 2018 CalumMechanicsElectric Field MeterEfm 002 he leak test

A) Start pump down 4.1e-10 torr l/s

after 1 minute 1.5e-9 torr l/s

B) vent and pump down while still in bag with he

this time levels out at 5.5e-10 torr l/s within 1 minute p1=7.5e-3 torr 

 

  61   Tue May 22 12:40:54 2018 CalumMechanicsElectric Field MeterEfm 0003 he leak test #2

EFM 003 2nd try with he leak test

2a) pump down and then fill bag

start 1.4e-10 mbar l/s

after 1 minute or so went up to 1.4 e-9 mbar l/s

2b) stopped, vented and then pumped down with he still in bag - also changed units to torr l/s

after 30 seconds  or so flat out at around 1.2e-9 torr l/s, was still at this value after 5 minutes

Now have really good system for leak testing with new bags, tapes, hoses out window etc 

 

 

  60   Mon May 21 15:49:04 2018 Rich AbbottGeneralElectric Field MeterEFM 2 and EFM 3 Update

Luis, Calum, Rich

Changed out the resistors in EFM 2 from 249k to 49.9k (R15), 240k to 43k (R12), 100pF to 470pF (C5 and C8) to allow wider CMRR range and reduced gain.

Bagged EFM 3 for helium leak test

20 sec helium. Pump rate started at 2e-10 mbar liter/sec helium and rose to 2e-9 over a few minutes.  Pressure then started slowly dropping to 1.9e-9.  We consulted with Dennis on this and he feels the result is marginal.  We elected to remove the unit from the bag and start again.  We checked the torque on each bolt using the same torque wrench as we used on EFM 1 in case there was a calibration error on the torque wrench.  Some bolts did move by 5 to 10 degrees or so.  The feeling we got was that the observed pressure rise may have been due to one of the couplings used to attach the leak checker to the EFM, so we bagged all the couplings as a precaution.  We will wait until the pressure is back down in the minus 10 range and try again.

We tried again, but are still having problems.  The leak rate is now ~7e-9 so we abandoned the helium leak check of EFM 3 and are moving to do the electrical tests of EFM 2.

Below can be seen the video from EFM3 test process.

 

 

Attachment 1: EFM_3_Leak_Test___Fail.mp4
Attachment 2: EFM_3_Leak_Test__Take_2.mp4
Attachment 3: EFM_3_Leak_Test_Take_3.mp4
  59   Fri May 18 16:30:29 2018 Rich AbbottElectronicsElectric Field MeterElectrical tests complete on EFM 3

Nichole, Calum, Rich

The electrical testing for EFM 3 is complete.  EFM 3 has now been torqued down on 5 out of 6 flanges to 180 inch-pounds and was left attached to the helium leak checker.  The leak checker was running well around 8e-10 mBar-l/sec as it was left.  A new procedure (E1800151) has been written to describe the following measurements.  The data below was used to seed the procedure with reasonable numbers.  A SPICE simulation was performed to validate the anticipated magnitude and phase plus noise performance.

Testing EFM3 Continued

Calculated RC pole frequency based on R=49.9k, C=470pF to be 6.79kHz. Using this as a basis for TF acceptance testing. X and Y pots set to 0561 HEX at start of measurement to get +/- gains to be nominally the same. SR-785 on 400 points, 100Hz to 10kHz, 1V source.

In on X+ -> Out Differential -> 100Hz, 6.024dB, -0.9 deg

In on X+ -> Out Differential -> 6.754kHz, 2.979dB, -45.3 deg

In on X- -> Out Differential -> 100Hz, 6.020dB, 179.1 deg

In on X- -> Out Differential -> 6.754kHz, 2.921dB, 134.3 deg

In on Y+ -> Out Differential -> 100Hz, 6.021dB, -0.8 deg

In on Y+ -> Out Differential -> 6.754kHz, 3.278dB, -43.4 deg

In on Y- -> Out Differential -> 100Hz, 6.021dB, 179.2 deg

In on Y- -> Out Differential -> 6.754kHz, 3.229dB, 136.5 deg

Next, the noise measured differentially at the output of each axis is measured with both positive and negative inputs shorted

X-axis noise at 100Hz -> 215 nVrms/rtHz

Y-axis noise at 100Hz -> 215 nVrms/rtHz

Measured the time constant of each input:

X+ -> 5.8 sec

X- -> 6.2 sec

Y+ -> 6.0 sec

Y- -> 5.8 sec

​CMRR (added by Luis on 052418)

X = -65dB @ 10Hz

Y = -75dB @ 10Hz

  58   Thu May 17 17:14:18 2018 Rich AbbottElectronicsElectric Field MeterInitial Test of EFM3

Luis, Rich

The attached notes are a first look at EFM3 after Luis soldered the new parts (C5 and C8 were 100pF and now 470pF for both X and Y differential amplifiers.  The resistors (R12 and R15) were 240k and 249k respectively, and are now 43k and 49.9k respectively. The resistor change was done to broaden the range of gain mismatch that we could compensate with the CMRR trim potentiometer. We measured the DC offsets with the inputs shorted and the DC response to a 1VDC input on all axes

Attachment 1: Testing_EFM3.pdf
Testing_EFM3.pdf
  57   Thu May 17 15:24:07 2018 Luis SanchezGeneralElectric Field MeterEFM No 1 Shipment to LHO

Stephen, Luis,Liz.

We bag the EFM this morning and Liz Help us to shipped the Electrometer to Hanford.

Attachment 1: EFM1.jpg
EFM1.jpg
  56   Wed May 16 14:39:28 2018 Rich AbbottElectronicsElectric Field MeterSetting Common Mode Rejection Ratio on EFM1

Luis, Rich

Today we explored setting the common mode rejection ratio (CMRR) of the Y-axis on EFM1.  We quickly realized we have a shortcoming in the range of CMRR we can adjust.  Based on the schematic used in EFM1 for the differential amplifier, we are only able to vary the gain (which in turn specifies the window aperture of gain that we can accommodate) over a 0.11dB range.  The intrinsic imbalance of a pair of inputs is usually around 0.3dB or so, so this window is not large enough to reach a minimum in the CMRR.  We are going to ship this unit as it is, but are immediately going to fix the problem and ready EFM2 as a replacement at LHO.

Attachment 1: EFM_Schematic_For_CMRR.pdf
EFM_Schematic_For_CMRR.pdf
  55   Tue May 15 19:47:01 2018 Rich AbbottElectronicsElectric Field MeterElectrical Tests on EFM 1, 2, and 3

Luis, Stephen, Calum, Liz, Rich

Today we had a first look at EFM 1, 2, and 3 in terms of the DC transfer function (observing 1VDC injected into each axis input results in 10V differential at the output), and the exponential decay associated with a step voltage change applied to a single calibration plate.  We spent the first part of the day in a typical state of confusion as we struggled to interpret the behavior of EFM 2 and 3.  EFM 1 was just clearing the final clean and bake process in the 40m Lab.  EFM 2 and 3 appear to be functional, however the background electrical noise in the Downs 2nd floor clean room seemed to be getting rectified and contributed a current to the input that lead to saturation.  At some point, we got the word that the mission critical EFM 1 was on it's way over from the 40m Lab, so we abandoned EFM 2 and 3 for now.

On EFM 1, we first verified all axes are functional with a simple DC injection.  1V injected into each output produced the requisite 10V differentially at the EFM outputs.  Then we moved on to the exponential decay test.  We first estimated the capacitance of each of the components on a single axis input to the EFM.  Corroborated by direct measurement where possible, we calculated based on area the following (1 inch PEEK standoffs on the sense plate to the body of the EFM, 1/2 inch PEEK standoffs from the sense plate to the calibration plate:

Sense Plate to EFM Body - 3.6pF

Sense Plate to Grounded Calibration Plate - 5.7pF

Ceramic Vacuum Feedthrough (measured) - 4pF

This yields a total expected capacitance of around 14pF including a bit for the input to the opamp internal to the EFM.

Well, so much for that notion.  Nature had a different picture of the capacitance.  As can be seen in the attached decay plot taken with an oscilloscope in response to a 5V step applied to the calibration plate, the RC time constant is around 23 seconds.  Assuming we got what we paid for in terms of the 10% tolerance, 10^12 ohm resistors, that would imply a capacitance of around 23pF.  Quite a bit more than we expected.  We then embarked on a process to try to figure out where we went wrong.  We had been shielding the sense and calibration plates during the decay measurement by creating a grounded aluminum foil tube that acted as an electric field shield.  We wondered if this was contributing and parasitic capacitance, so we repeated the decay with no shield.  Now we had a 19.8 second RC time constant.  Not much different, but better.  Next, we hung the sense and calibration plates off the side of the work bench in a effort to see if the proximity to the grounded bench was a factor.  Now we had a 18.8 second RC time constant.  Next we removed the voltage source from the calibration plate to effectively unground it so it would be an unreferenced metal object that theoretically should add no capacitance. This also required that we make a big step change in the local EFM electric field environment (Rich wiggling around) such that it would essentially saturate as we no longer had a calibration plate to drive.  This brought the time constant down to 13.6 seconds.  Lastly we did the absurd and completely removed the calibration plate even though Rich was sure it would make absolutely no difference.  Now the time constant decreased to 8.8 seconds to his utter amazement.  Here's a summary:

RC Time Summary
Note RC Time Constant (sec)
Grounded foil Shield around sense and cal plate 20.4
Foil shield removed 19.8
Plates hanging off bench 18.8
Ungrounded Calibration Plate 13.6
Calibration Plate Removed 8.8

 

 

 

 

 

 

Perhaps, and most likely by coincidence, the time constant of 8.8 seconds actually is beginning to agree reasonably with the anticpated capacitance of the sense plate plus feedthough (3.6pF plus 4pF)

Attachment 1: EFM1TimeConstXminusAxis.jpg
EFM1TimeConstXminusAxis.jpg
  54   Mon May 14 13:31:59 2018 Luis SanchezGeneralElectric Field MeterEFM Cube No. 1 Bake at 50C for 24hrs

Calum, Rich, Bob, Luis

EFM Cube No.1 was loaded into oven F for a 24hr 50C bake. The bake started around 11:00AM today 051418.

Attachment 1: Cube_No1_24hrs_50C_bake.jpg
Cube_No1_24hrs_50C_bake.jpg
Attachment 2: Cube_No_24hrs_50C_bake_2.jpg
Cube_No_24hrs_50C_bake_2.jpg
  53   Mon May 14 10:57:43 2018 Stephen AppertProgressElectric Field MeterMechanical Part History and Serial Numbers

WIP (photos will be posted, more information will be gathered)

  1. 14 May 2018: Stephen and Luis gathered serial numbers from Cube 2 / Cube 3 / spares (all of which were Class A baked in Bake-9049) (ICS records will be updated)
    1. Cube 3
      1. D1800043-v1 SN003
      2. D1800032-v2 SN013
      3. D1800032-v2 SN004
      4. D1800032-v2 SN001
      5. D1800032-v2 SN012
      6. D1800054-v1 SN002
      7. Cube 408003 unserialized
    2. Cube 2
      1. D1800043-v1 SN004
      2. D1800032-v2 SN009
      3. D1800032-v2 SN010
      4. D1800032-v2 SN002
      5. D1800032-v2 SN014
      6. D1800054-v1 SN003
      7. Cube 408003 unserialized
    3. Spares
      1. D1800032-v2 SN011
      2. D1800032-v2 SN007
  52   Fri May 11 20:37:10 2018 CalumMechanicsElectric Field MeterAccumulation test

Accumulation test data 

https://dcc.ligo.org/LIGO-T1800212

 

Attachment 1: IMG_0007.JPG
IMG_0007.JPG
  51   Fri May 11 11:56:32 2018 Luis SanchezGeneralElectric Field MeterStill chasing the 1teraohm resistor

Luis, Rich,

Yesterday we started a new test. This test was setup trying to confirm the resistance from the 1 tera ohm resistor. This time we have the idea of adding a known capacitor in parallel with the setup. Now we have 2 equations with 2 unknowns. 
1st equation is  R x C =6sec; the charging time of 6sec was found in the experiment developed yesterday. 2nd equation is R x (Csense + C) = charging time. Csense is the capacitance from the flange to the sense plate, 
we measured the separation of these plates and found it to be 0.656inch apart. With this data we found the capacitance of 5.4pF.  We measured the new  RC time and we got 24sec. 
With this new data we have the 2nd equation as  R(5.4pF + C)=24sec if we solve for C considering that R=6/C we found that C= 1.8pF making R =3.3teraohms. Having the time of 24sec makes the resistor larger than 1 tera ohm and also 
make us think that this new setup is adding more capacitance to our system. The way we applied the input signal to the sensor created a new capacitor? we need to measure the new capacitance!!!
It is important to mention that adding a sense plate to our original set up(without isolating the pcb boards)made a very good detector but the signal is too noisy, with this signal it was dificult to identify the asymptotic line when trying to fit the curve. 
We tried to isolate the sense plate for any external signals but the shield applied to setup was not enough, the signal behaved almost similar as before.     


Today we started with a good idea of adding the calibration plate to the set up (the separation between this plate and the sense plate is about 0.5inch). We shielded the set-up completely and after taking this new ring down measurement we got around 18pF which give us a time constant of 18.2seconds at the 63% (1.79v) charge time from the top down to asymptote. These measurement was considering the max peak voltage at 2.62v and the asymptote line at -0.22v. With this new piece of information we conclude that our Resistor is indeed a 1Teraohm resistor. 18.2seconds/18pF= 1.011x10E12ohms.See scope image. 

Attachment 1: Ring_down_sense_plate_not_shield.PNG
Ring_down_sense_plate_not_shield.PNG
Attachment 2: Cal_plate.jpg
Cal_plate.jpg
Attachment 3: Cal_plate_separation.jpg
Cal_plate_separation.jpg
Attachment 4: Sense_plate_separation.jpg
Sense_plate_separation.jpg
Attachment 5: Set_up_partially_shielded.jpg
Set_up_partially_shielded.jpg
Attachment 6: Set-up_shielded.jpg
Set-up_shielded.jpg
Attachment 7: Ring_Down_measurement_Electrometer.jpg
Ring_Down_measurement_Electrometer.jpg
  50   Thu May 10 10:19:01 2018 Rich AbbottElectronicsElectric Field MeterMeasurement of input capacitance of EFM

Luis, Rich

Yesterday, we spent quite sometime messing around trying to measure the input capacitance of the EFM.  Each EFM input feeds into an AD549 opamp.  These opamps have a very high input impedance made using JFETs.  Essentially, one is looking into the gates of a bunch of JFETs in parallel.  The gate of a JFET can be envisioned as a very high impedance resistor (>10^15 ohms) in parallel with a picofarad or so.  We thought we would be clever and measure this capacitor.  It turns out that the opamp had different ideas...

While the input of the opamp is reasonably modeled by a parallel resistor and capacitor, the capacitor value depends on the amount of voltage present on the opamp input.  When we would try to measure the input impedance using a simple capacitance meter, the meter would apply a sinusoidal voltage that is used to calculate the capacitance of the circuit hooked to the meter.  As the voltage of the sinusoid varies, so does the capacitance.  This dynamic situation confused the hell out of the capacitance meter, which staunchly refused to place any limit on the measured capacitance other than approximately a value from minus 1uF to plus 1uF.  The actual expected value should be around 1pF.

Eventually, we gave up trying to measure the input capacitance and instead applied a step change of input voltage to the electrometer input (using a piece of aluminum foil near the copper rod that forms the input to the EFM).  We would watch and plot the voltage decay as a function of time.  Because we don't know the resistance (assumed to be 1 x 10^12 ohms), nor do we know the input capacitance, we can only derive the product, RC, of the two.  Here's what we saw:

This grainy image shows some measured data with a fit overlayed in orange.  If the resistance is indeed 10^12 ohms, then the calculated capacitance would be 6pF.  We were able to measure the capacitance of the ceramic feedthrough to be about 4pF, so this means the total input capacitance of the EFM circuit board and opamp would be around 2pF.  Quite reasonable really.  Next, we are going to bolster our confidence that these resistors are actually producing 10^12 ohms by using a second resistor in series with the EFM input to form a 2:1 divider.  If this doesn't end tragically, we will post the results soon.

Attachment 1: RCtimeEFM.pdf
RCtimeEFM.pdf
  49   Thu May 10 09:45:11 2018 Luis SanchezGeneralElectric Field MeterHelium leak test video

Bob, Calum, Rich, Luis.

As Rich mentioned, the EFM helium leak test has passed.

I am attaching a video that was taken during this test, in this video we can see how the vacuum pressure was kept during the experiment eventhough the tracer gas was in the vecinity of the EFM.

DCC document E1800144 details some aspects from the test process and describe the final results.

Finally, Today (051118) I was able to compress the video file and this can be viewed here. 

Attachment 1: Helium_Leak_Test.mp4
Attachment 2: Helium_Leak_Test.mov
  48   Tue May 8 16:57:28 2018 Luis SanchezGeneralElectric Field Meterpre assemble EFM

I started to assemble the interface board power regulators and sense board in clean room located in dows 2n floor. EFM Cube2 and Cube3 are partially assemble, sense boards and interface board are connected thru the interface wires. One cable for each sense board.

 

Attachment 1: Shield_Plate.jpg
Shield_Plate.jpg
Attachment 2: Interface_board.jpg
Interface_board.jpg
Attachment 3: Interface_board_Flange.jpg
Interface_board_Flange.jpg
Attachment 4: Sense_boards.jpg
Sense_boards.jpg
  47   Mon May 7 17:53:46 2018 Rich AbbottElectronicsElectric Field MeterCapacitance Measurement

Luis, Rich

In an effort to understand where the ~20pF capacitance comes from as measured at the site of the sense plate to body, we measured the capacitance of the ceramic feedthrough all by itself.  After checking two examples, we concluded that the feedthrough capacitance is about 3pF.  The datasheet for the AD549 opamp says the amplifier input capacitance is around 1pF.  A calculation of the approximate capacitance of the sense plate to the body is around 7pF ignoring fringe fields, and the capacitance of the L-bracket inside the cube to the inner flange is around 0.5pF.  This total (~11.5pF) leaves 8.5pF unaccounted for.

  46   Mon May 7 15:06:19 2018 Luis SanchezGeneralElectric Field MeterEFM No1 Bake process at 50C in oven D C&B Caltech

Luis, Bob, Stephen, Calum,

The EFM No1 was loaded into oven D and the bake process started with set point of 50 Celcius, Nitrogen was used during the process. It is possible that the neon acumulation test will be happen tomorrow (050818).

The sense feedthrough pins were grounded using aluminum foil (which connects to all 4 pins).

In the mean time we will start to assemble the Power regulators on Interface board(D1800040) and set the brackets to connect the D15 feedthrough flange.

Also several items (related to EFM) were unload from oven F and move to float bench for overnight cooling down.

Attachment 1: Loading_EFM_into_oven_D.jpg
Loading_EFM_into_oven_D.jpg
Attachment 2: EFM_No1_inside_oven_D.jpg
EFM_No1_inside_oven_D.jpg
Attachment 3: EFM_in_Oven_D.jpg
EFM_in_Oven_D.jpg
  45   Mon May 7 10:36:50 2018 Rich AbbottGeneralElectric Field MeterMoving EFM #1

Rich, Bob, Calum, Todd, Luis

EFM#1 has passed the helium leak checking phase and is being moved over to the 40m Lab for the neon accumulation test.  The electrical inputs will be banded with clean aluminum foil to protect against inadvertent static electric charge damage.

  44   Fri May 4 15:58:20 2018 Rich AbbottGeneralElectric Field MeterNeon Gas Fill of Cube 1

Todd, Luis, Calum, Rich

Filled EFM #1 with neon today.  This is the first time we have done this so a procedure (E1800138) was written.  Basically we wrapped the EFM with a band of foil across all electrometer (X and Y) inputs to ground them, then put the EFM into a bag.  We used a vacuum cleaner to remove the air from the bag, then sealed off the port used for the vacuum cleaner.  A hose to inject neon from the gas bottle was prestaged such that it went into the top flange port of the EFM (the blank off flange).  We filled the bag with neon while recording time and flow rate.  The neon was being delivered at ~4psi which corresponded to ~1 +/-0.1 standard cubic feet per minute as measured by a sightglass flow meter.  It took 1 minute and 50 seconds to fill our bag.  Using flow and time this corresponded to 1.83 cubic feet of neon.  We also measured the dimensions of the bag which had formed a rough cylinder (2.5ft long, 0.5ft radius) which equates to 2 cubic feet which is in good agreement with the flow based number.

After filling the bag with neon, we found that it was better to snug up the bolts using a closed end wrench thus keeping the neon inside the EFM, then to do the final torquing immediately after the EFM is removed from the bag.  The torque spec used was 180 inch-pounds to bolt the CF flange to the cube.  This spec was derived from the MDC vacuum specification for the cube and flange arrangement we used.

Attachment 1: Under_vacuum.jpg
Under_vacuum.jpg
Attachment 2: Sealed_bag.jpg
Sealed_bag.jpg
Attachment 3: 180_inch-pounds_torque.jpg
180_inch-pounds_torque.jpg
Attachment 4: Bolted_after_fill_with_Neon.jpg
Bolted_after_fill_with_Neon.jpg
Attachment 5: TeamWork.jpg
TeamWork.jpg
Attachment 6: TeamWork2.jpg
TeamWork2.jpg
Attachment 7: EFM1.jpg
EFM1.jpg
Attachment 8: NeonBottle.jpg
NeonBottle.jpg
Attachment 9: NeonInflatedBag.jpg
NeonInflatedBag.jpg
  43   Fri May 12 16:34:54 2017 AlenaProgressLaser DamageFused silica viewport laser damage test

Made series of laser irradiation on a 3” fused silica uncoated optic https://dcc.ligo.org/LIGO-S1700118-v4

The optic is planned to be used as a viewport at LDF. In air laser damage test is required before using the optic as a viewport on the LDF vacuum chamber. Beam on the target is 127 by 91 microns ellipse. See estimated equivalent laser power input table and layout picture in the related doc https://dcc.ligo.org/LIGO-T1700184-v1

No damage was observed using DF microscope. The inspection was done using manual control – no scanning. After each irradiation run a picture was taken with a FLIR camera.  Post irradiation pictures taken with DF microscope are posted herte https://dcc.ligo.org/LIGO-S1700118. 

Warming up of the beam dumps up to 57 deg C was detected. No other elements of the set up were changing temperature (including the target – 3” fused silica optic and it`s mount)

  42   Wed May 3 17:20:24 2017 AlenaProgressLaser DamageFused silica viewport laser damage test

Added a magnetic base for the 3" diameter target mount. The mount is angled in order to dump the reflection from uncoated fused silica target. Beam on the target is 127 by 91micron ellipse. Irradiation is planned at 25, 50, 75 and 100% power (100%=50W). Equivalent power densities and irradiation runs see in the dcc doc https://dcc.ligo.org/S1700118-v1

Attachment 1: unnamed.jpg
unnamed.jpg
  41   Mon May 1 17:14:41 2017 AlenaProgressLaser DamageFused silica viewport laser damage test

Worked on a laser layout for in-air fused silica viewport optic laser damage test. Made a temporary enclosure to prevent high power laser scatter and to stop any possible fragments of potentially damaged viewport. A “labyrinth” shape was built instead of making holes for the laser tube. The top of the enclosure will be covered with one large black panel as well. The beam was focused to 2w=100 microns at the position of the target (red line)

Attachment 1: 20170501_163511.jpg
20170501_163511.jpg
  40   Thu Apr 27 15:08:21 2017 AlenaOpticsInspectionelliptical mirrors batch 2 (thin mirrors)

Inspected batch 2 of thinner elliptical mirror received from Brand Laser Optics & Mfg coated on one side. Uncoated surface is matte vs glassy coated side. Shiny green reflections can be observed on the coated side. The mirrors have no chipping but a bit dusty from the package. I have marked the front and back side of the mirrors on the box.

  39   Tue Apr 25 14:14:14 2017 AlenaOpticsInspection 

Elliptical mirror received from Brand Laser Optics & Mfg is coated on one side. Uncoated surface is matte vs glassy coated side. Shiny green reflections can be observed on the coated side. I have also confirmed it using a 1064 laser and a laser card. The mirror has chipping at least on one edge. Surface looks dusty from the packaging but clean. I have marked the front and back side on the box.

  38   Wed Apr 5 13:27:30 2017 AlenaProgressPZT jitter experimentD1700002 v-block test

New mount has been reworked  https://dcc.ligo.org/D1700002 in order to use 2 v-lamps. See modefied mount https://dcc.ligo.org/LIGO-D1700002-v5

See T1600060 the transfer function measured with four different configurations: - elliptical mirror, no viton (black curve)
- elliptical mirror, viton under tip of the PZT (blue curve)
- elliptical mirror, viton (full length) (green curve)
- 2" mirror, viton (full length) (red curve)

First resonance appears at about 2 kHz which is very close to the internal resonance frequency of the PZT (3 kHz with no load according to the specs). Adding viton in the grove did not dump the resonance (black curve vs green and blue) however it can be an option. Changing the mirror from 2" to a lighter elliptical is still a significant improvement even with the new mount (red curve vs others).

 

  37   Wed Apr 5 13:12:57 2017 AlenaProgressPZT jitter experimentRed v-block test

PZT`s transfer function measurement with a test v-block was done in tree differen configurations:

- no viton (blue curve)

- viton under PZT (orange curve)

- viton only under the front part of the PZT (green curve)

Conclusion: 3rd configuration is the winner. It provides stiff clamping of the back of the PZT plus dumping. Need to design new v-block type mount with two clamps. Promissing first resonance at 3kHz with the new mount

 

  36   Tue Jan 24 15:47:03 2017 AlenaGeneralPZT jitter experimentTape mount vs sleeve mount

PZT body was taped to a v-block. An extra mirror used in the layout for stirring the beam. Pictures show a comparison between this tape mount and a sleeve mount in a U200 mount. Taping the PZT damped the two resonances (300 Hz and 1 kHz). The internal PZT resonance was stiffen.

Attachment 4: Ppr.png
Ppr.png
  35   Thu Jan 12 07:36:55 2017 StephenProgressCASI BRDFGigE Camera: Photos of Oxidized SSTL

This log entry reflects on recent efforts to image the 1064nm spot on samples undergoing BRDF measurement within the CIT CASI scatterometer. The cameras used are (reiterating from eLOG ENG_Labs/33)

  • Basler ace acA640-100gm, same as used at sites, borrowed from Johannes at 40 m. Equivalent to acA640-120gm, which is the current Basler designation of this camera.
  • Basler scout scA1400-17gc, originally purchased by OpLev design team, leftover from aLIGO efforts, borrowed from Eric/Gabrielle in West Bridge.

The set up and measurement is described by the following procedure:

  1. A sample of oxidized stainless steel was found in the lab and taped to an existing sample mount using kapton tape.
  2. Images were taken with the room lights on and then without the room lights.
  3. Imaging settings were maintained for each photo taken with an individual camera, but gain values and packet size varied between cameras.

Hardware for the cameras included:

  • a lens with adjustable focus (Rainbow, 50mm focal length, C mount, 2/3" to 1" aperture, 1:1.8 image; 40m loan)
  • where indicated, a ~70% transmissive bandpass filter (Newport 10LF25-1064).
  • mounting post with fixed clamping platform (will add photo), approximately 1 meter from the sample.

The acquired images may be summarized by the following points:

  • The 1064nm light was represented by the color scout camera as white-ish pixels with a bluish tinge.
  • The scout camera appeared to have a number of dead pixels (someone with more camera expertise may have better terminology; I am referring to the scattered red and blue pixels that do not appear to be real light).
  • The focal length of the two cameras was different, hence the difference in perceived spot size.
  • The bandpass filter reduced intensity of the image and, with the light off, did not seem to provide any useful noise cancelling.

Image 07: Lights on, scout OpLev GigE Camera Link to casi_test_07_laptop_lights_high_exposure_10_mw_high_gain.bmp

Image 08: Lights off, scout OpLev GigE Camera Link to casi_test_08_no_lights_high_exposure_10_mw_high_gain.bmp

Image 09 Lights off, bandpass filter, scout OpLev GigE Camera Link to casi_test_09_no_lights_high_exposure_10_mw_high_gain_bandpass_filter.bmp

Image 13: Lights on, ace Site GigE Camera Link to casi_test_13_site_gige_foyer_light_auto_exposure_10_mw_auto_gain.bmp

Image 15: Lights off, ace Site GigE Camera Link to casi_test_15_site_gige_foyer_light_auto_exposure_10_mw_100_gain.bmp

Image 19: Lights off, bandpass filter, ace Site GigE Camera Link to casi_test_19_site_gige_no_light_auto_exposure_10_mw_100_gain_bandpass_filter.bmp

Attachment 1: casi_test_07_laptop_lights_high_exposure_10_mw_high_gain.bmp
Attachment 2: casi_test_08_no_lights_high_exposure_10_mw_high_gain.bmp
Attachment 3: casi_test_13_site_gige_foyer_light_auto_exposure_10_mw_auto_gain.bmp
Attachment 4: casi_test_15_site_gige_foyer_light_auto_exposure_10_mw_100_gain.bmp
Attachment 5: casi_test_09_no_lights_high_exposure_10_mw_high_gain_bandpass_filter.bmp
Attachment 6: casi_test_19_site_gige_no_light_auto_exposure_10_mw_100_gain_bandpass_filter.bmp
  34   Wed Jan 4 14:08:46 2017 AlenaProgressLaser DamageFused silica viewport

Measured lasaer power at the end of the layout with defould pyrex viewport, fused silica AR coated viewport and no viewport. See https://dcc.ligo.org/T1700003-v1 for more info.

Laser power measured using a “PM100USB” power meter and an S314C sensor (±3% measurement uncertainty at 1064 nm). Gray dots represent the power indicated at the laser display which can be inaccurate. There are additional mirror losses at the beam transport via RTS setup. That is why power at LDF setup location will differ from the values shown on the laser display.

Black curve on figure1 is the power measured at the end of the LDF laser layout (including a window from the viewport cover) with no viewport (figure 2). Blue and red curves represent the power measured with a fused silica and pyrex glass correspondently. In both cases the glass was added at the end of the laser layout (see figure 3).

Figure 1: 1064 laser power measured as a function of the input value

Figure 2: Power-meter installed at the end of the LDF laser layout (correspond to the black curve on figure 1)

Figure 3: Power-meter installed at the end of the LDF laser layout plus the fused silica glass (correspond to the blue curve on figure 1)

Almost 100% transmission trough the fused silica viewport has been observed. Losses in fused silica were less that the resolution of the power meter. In contrast, losses up to about 15% were observed in the pyrex glass. An integrating sphere may be used for more accurate measurement.

Calibration function between the display input and the power in vacuum chamber after passing the viewport (±3% measurement uncertainty in not considered):

PLDF[W]=-4.796812+ 0.549683 Pdisplay input [%]

 

 

 

 

  33   Thu Dec 1 18:12:17 2016 StephenHow ToCASI BRDFGigE Camera: First Set-Up

Summary

I was successful in connecting two Basler GigE cameras to my personal laptop and viewing/saving images using pylon, the UI downloaded from the Basler website here. The details of the cameras are documented below:

  • Basler ace acA640-100gm, same as used at sites, borrowed from Johannes at 40 m. Equivalent to acA640-120gm, which is the current Basler designation of this camera. ~$500 for qty of 1, 3 week lead time.
  • Basler scout scA1400-17gc, originally purchased by OpLev design team, leftover from aLIGO efforts, borrowed from Eric/Gabrielle in West Bridge. ~$3300 for qty of 1, 3-4 week lead time.

There will be more to follow in subsequent log entries regarding integration into the CASI BRDF measurement system.

Some Helpful Resources

aLIGO GigE camera layout and usage. Joe Betzwieser from LLO has a guide to the camera usage and system set-up at the sites located at T1300202.

Basler set-up guide. Basler has a very good set up guide which covers details and some troubleshooting. Accessed at this link, by navigating to [ Support > Downloads > Current pylon Windows ]

Description of Set-Up

Once all of the materials were in hand, a PoE injector (TP-Link TL-POE150S; 40m loan) was connected in series between my laptop and the camera in question. I also had success using the power supply supplied with the camera and a direct ethernet connection to the laptop. A lens (Rainbow, 50mm focal length, C mount, 2/3" to 1" aperture, 1:1.8 image; 40m loan)  was installed on the camera, and the focus ring was adjusted as needed (no iris adjustment required).

The pylon IP Configurator was used to force IP addresses at the onset (Static IP used for the ace and DHCP used for the scout, based on defaults upon opening). The camera status was "OK", then the pylon Viewer was opened, the camera was opened (may need to refresh the list) and placed in Continuous Mode for image collection (ctrl-S to save an image) See the Basler set-up guide (link at the top of this post) for full details of camera set-up.

Difficulty with Set-Up

The only real difficulty I had turned out to be due to exceeding the bandwidth of the set-up. The common error was:

Error: 0xe1000014 "The buffer was incompletely grabbed. This can be caused by performance problems of the network hardware used, i.e. network adapter, switch, or ethernet cable. To fix this, try increasing the camera's Inter-Packet Delay in the Transport Layer category to reduce the required bandwidth, and adjust the camera's Packet Size setting to the highest supported frame size."

Accessing Transport Layer in the camera drop-down and following the recommendations eliminated the bandwidth issues. I was confused at first because there is a second Transport Layer category that lives outside of the camera drop-down. This one is not useful.

Better yet, following Joe's recommendations in "Notes on camera settings" on page 11 of T1300202 (link at the top of this post) made for a seamless camera run.

Results

The color image from the scout camera was taken from about 3m distance with my unsteady hand holding the camera.

Link to scA1400-17gc_image_from_test_office_emergency_map_01.bmp

The black and white images from the ace camera were taken from the same 3m distance, also mounted in an unsteady hand. The image tagged 02 has an increased Gain (250 --> 700) which increased the contrast while causing the image to be less sharp and detailed.

Link to acA640-100gm_image_from_test_office_light_switch_01.bmp

Link to acA640-100gm_image_from_test_office_light_switch_02.bmp

Admittedly, these images don't really mean anything. There will be more to follow in subsequent log entries regarding integration into the CASI BRDF measurement system.

 

 

Attachment 1: scA1400-17gc_image_from_test_office_emergency_map_01.bmp
Attachment 2: acA640-100gm_image_from_test_office_light_switch_01.bmp
Attachment 3: acA640-100gm_image_from_test_office_light_switch_02.bmp
Attachment 4: scA1400-17gc_image_from_test_office_emergency_map_01.bmp
Attachment 5: acA640-100gm_image_from_test_office_light_switch_01.bmp
Attachment 6: acA640-100gm_image_from_test_office_light_switch_02.bmp
  32   Thu Dec 1 15:14:57 2016 AlenaElectronicsPZT jitter experimentrepeatability test

Two front screws on the PZT sleeve mount out and back in to checj the repeatability. No big difference has been observed.

Attachment 1: 20161130_155208.jpg
20161130_155208.jpg
  31   Wed Nov 30 15:49:48 2016 AlenaElectronicsPZT jitter experiment 

I measured the transfer function of the PI`s PZT with two configurations of the setscrews that hold the PZT`s body in the sleeve mount: 4 screws and only 2 back screws. The natural frequency resonance slightly moved to the lower frequency range with 2 screws setup. Also there is a bit more mess in the high frequency range with the two screws configuration. But it is hard to see clear in this range using the current setup.

 

 

 

  30   Tue Nov 29 13:45:13 2016 AlenaGeneralThomasSR3 actuator test

The temperature was measured by two thermocouples: integrated thermocouple at the ring heater and external thermocouple at the back side of the radiative plate (gold plated). Four data sets were taken (with no gasket - gray,  with indium gasket - green, tin -indium alloy gasket - blue and tin-indium "heat spring" - magenta on the graph). The higher temperature is the ring heater and the lower temperature is the external thermocouple at the plate. The two temperatures are different due to the heat losses caused by low thermal contact between the ring heater and the radiative plate. For example, when the ring heater at 120 deg, the radiative plate is only about at 75 deg C. Using a gasket prevents overheating of the ring heater. Tin-indium gasket reduces the heater temperature to 90 deg C, indium - to 80. Tin-indium has lower thermal resistance than indium by is nor as soft as indium therefore indium performs better. Making a tin-indium "heat-spring" improves the contact such that it performs as good as pure indium, at the same time it is easier to handle (not too soft). But tin-indium melting temperature is 117 deg C and pure indium melting temperature is 156 deg C. Therefore indium is the winner.

 

Fig1. Temperature measured at the ring heater and radiative plate with elevated current (200 mA, 250 mA, 300 mA)

 

 

Fig2. In heat-spring gasket installed between the heater and the plate

 

 

Fig3. InSn gasket installed between the heater and the plate

 

 

Fig4. InSn heat-spring gasket

Attachment 1: gaskets_test.png
gaskets_test.png
  29   Thu Nov 10 16:12:23 2016 AlenaGeneralPZT jitter experiment2" mirror vs elliptical 25 mm minor axis mirror

Swapped 2" mirror with an elliptical mirror (25mm minor axis)

  28   Tue Nov 8 16:23:58 2016 AlenaElectronicsPZT jitter experiment1" miror vs 2" mirror

setup i  - 1" mirror

setup ii - 2" mirror

I have swapped 1" mirror (5 mm thick) with a 2" mirror (12 mm thick). I just remover the old mirror and glued the new one in place without moving od disassembling the PZT sleeve mount and u200 mount. the resonance at about 300 Hz came back particularly in Y and a new resonance appeared at about 800 Hz in both: Pitch and Yaw, X and Y. This is probably the PZT`s natural frequency resonance. The specs mentioned 1.6 Hz  resonance under the load of a 25 mm (8 mm thick) mirror.

 

 

  27   Tue Nov 8 16:09:12 2016 AlenaElectronicsPZT jitter experimentSetup 6 vs i vs ii

Setup 6 - 1" mirror, magnetic base , glue and so on (final setup used for presentation )

Setup i change - back to the orange base

Setup ii change - Ø2" mirror 12.5 mm thick

Isolating the table did not change the results at all

 

  26   Wed Oct 19 14:22:11 2016 AlenaElectronicsPZT jitter experimentsetup 6

More epoxy!.. Nor much improvement byt the remaning peak moves. So that 300 Hz is related only to the connection of the two mounts.

Attachment 3: _TF6_Y_.png
_TF6_Y_.png
Attachment 4: setup6.jpg
setup6.jpg
  25   Wed Oct 19 12:40:21 2016 AlenaElectronicsPZT jitter experimentsetup 5

After halfway success with a drop of epoxy between the two mounts  I took it apart again, cleaned the dry epoxy and put really a lot of fresh epoxy. This removed "pitch" and "yaw" resonances on X, however small peaks are still observed on Y.

Attachment 3: setup_5_pzt.jpg
setup_5_pzt.jpg
Attachment 4: setup5.jpg
setup5.jpg
Attachment 5: Yaw_TF5_.png
Yaw_TF5_.png
Attachment 6: Pitch_TF5_.png
Pitch_TF5_.png
  24   Wed Oct 19 12:33:39 2016 AlenaElectronicsPZT jitter experimentsetup 4

The PZT mount holds in the Newport mount with only one set screw. I released the screw and put a drop of 5 min epoxy between the two mounts. After this the "pitch" resonance was almost gone and the "yaw" resonance reduced.

  23   Tue Oct 18 16:40:11 2016 AlenaElectronicsPZT jitter experimentsetup 3

I glued the PZT mount to a magnetic post to see is it affects the resonance. The peek at about 300 Hz moved right and reduced the magnitude in Pitch

 

 

 

Attachment 2: setup_3.jpg
setup_3.jpg
Attachment 3: setup_3_pzt.jpg
setup_3_pzt.jpg
Attachment 5: Pitch_TF3_.png
Pitch_TF3_.png
Attachment 8: Yaw_TF3_.png
Yaw_TF3_.png
  22   Fri Oct 14 17:12:15 2016 AlenaElectronicsPZT jitter experimentsetup 2

I change setup to increase the distance between the QPD and PZT because I wanted to repeat the measurement at 0.1 V amplitude on the PZT. The new measurement showed that the reason of the bad curve was not the QPD accuracy. So probably the PZT riches it`s accuracy below 1 V. Also the resonance at about 300 Hz moved. I have tighten the PZT mount to the orange post differently this time (stronger). This could be a reason.

Attachment 1: Yaw_TF_Amplitude_0.1_set2.png
Yaw_TF_Amplitude_0.1_set2.png
Attachment 6: 20161014_170007_resized.jpg
20161014_170007_resized.jpg
Attachment 7: Pitch_TF_Amplitude_1_set2.png
Pitch_TF_Amplitude_1_set2.png
Attachment 8: Yaw_TF_Amplitude_1_set2.png
Yaw_TF_Amplitude_1_set2.png
Attachment 9: Yaw_TF_Amplitude_0.1_set2.png
Yaw_TF_Amplitude_0.1_set2.png
  21   Thu Oct 13 16:06:34 2016 StephenSummaryPMC WorkTesting of Spring Plate for PMC Spacer Assy

Abstract

The purpose of this test was to evaluate a spring plate in a custom optic mount design which will be deployed in the reworked and upgraded PMC Spacer Assy per D1600270.

Summary of Results (Fit)

Mechanical inspection of the fabricated showed conformance to manufacturing specifications. Assembly of the Fit Test Interface Assy D1600398 in the nominal configuration was successful.

  • The primary source of assembly clearance was the 0.030” radial gap between the mirror and the 1in. Mirror MT.
  • Secondarily, assembly clearance was evident in the clearance bore bolt circle in the 1in. Mirror MT. for the M2.5 x 0.45 SHCS.

Some additional consideration (Tooling? Interfacing part?) may be necessary to combat these sources of “slop” and permit repeatable assembly.

Summary of Results (Holding Force)

The Fit Test Interface Assy D1600398 was laid flat on the work bench with the normal axis of the mirror vertical. The force required to overcome the static friction for a given spring plate defection and induce macroscopic translation was evaluated as follows:

  • 0.050” deflection (nominal): 6 +/- 0.5 ounces of force (1.7 newtons)
  • 0.030” deflection: 4 +/- 0.5 ounces of force (1.1 newtons)
  • No spring plate: no measurement, as this force was too small for this experimental set-up

In comparison, the weight of the optic is estimated to contribute a force of 1.6E-2 newtons in the nominal orientation. The optic did not sag in the mount when held in the nominal configuration and orientation for 2 hours.

In shipping configuration, the set screws were found to couple rigidly to the barrel of the optic and provide sufficient static friction to withstand jarring motions. The design seems to be sound from a shipping point of view.

Summary of Results (Comparison with COTS)

Two Commercial Off-The-Shelf (COTS) optic mounts were investigated. Barrel contact points, namely two flats and a preloading point-contact plunger, were used in conjunction with a spring plate. The assembly was rigidly coupled with high static friction constraining all degrees of freedom.

Perhaps barrel contact points should be employed in the LIGO design.

 

-----------------------------------------------

Full Description of Experiment

-----------------------------------------------

 

COTS Spring Plate Mounts

Two models of Commercial Off-The-Shelf (COTS) Low Outgassing optic mounts from Newport which provide axial clamping using the COTS 906919-02 Mirror Mount Spring Plate were investigated. Both use the Spring Plate in conjunction with barrel contact points, namely two flats and a preloading point-contact plunger.

An optic is loaded into these mounts by using the barrel to depress the plunger, then tilting the optic into flush face contact with a retaining lip that is integral to the mount. The plunger then pushes the optic into its seat on the two flats (visible contacting the right hand site in the image below). Subsequently, the blade spring is bolted to the exposed face, and the assembly is rigidly coupled with high static friction constraining all degrees of freedom.

Optic Mount models are:

  • NewFocus (Newport) 902817-04 (9817-6-NI-K) (pictured below, with dirty optic)
  • NewFocus (Newport) 902816-04 (9814-6-NI-K)

 

Inspection

The parts procured for the Fit Test Interface Assy D1600398-v1 were received in good condition. Using micrometers, calipers, and pin gages, Stephen employed partial inspection of mating and critical features to establish that the parts were made to specification by the vendor. The parts consisted of the COTS 906919-02 Mirror Mount Spring Plate (Qty 25, New Focus / Newport), the D1600233-v1 1in. Mirror MT. (Qty 1, ProtoLabs), the D1600397-v1 Base Plate (Qty 1, ProtoLabs) and assorted COTS hardware and fasteners.

The only item worth raising from this inspection is the difficulty of inspecting the 1.010” Diameter pocket in the D1600397-v1 base plate. No accurate measurement could be made due to the nominal .030” high shoulder, the nominal .0075” machine tool radius, and the un-deburred lip, all of which limited the contact area and precision for the calipers. A fit check of the PZT showed that the pocket was large enough to insert the PZT, and this was taken to be sufficient to “Pass” the size dimension on this feature.

Assembly

To assemble the Fit Test Interface Assy, the following procedure was followed:

  1. The COTS 906919-02 Mirror Mount Spring Plate was bolted to the D1600233-v1 1in. Mirror MT. and the subassembly was placed flat on the workbench with the spring plate on the table side.
  2. The 25 mm Diameter x 5 mm Thick mirror was laid flat on top of the spring plate within the 1in. Mirror Mount, with the HR-coated side facing away from the table. Now, the AR-coated side was in contact with the Spring Plate.
  3. The M2 x 0.4 set screws were tightened onto the barrel of the mirror until rigidly coupled, with the mirror centered by eye in the 1in. Mirror Mount. THIS WAS NOT A VERY PRECISE CENTERING OPERATION, but using a pin gage as a spacer, I was able to position the optic within 0.005” of the center of the mount.
  4. The 1in. Mirror Mount was lowered toward the surface of the D1600397-v1 Base Plate, until the HR-coated side was registering on a single plane with three locations of point contact.

  1.  The M2.5 x 0.45 SHCS were tightened by a couple of turns, indexing the 1in. Mirror Mount BUT NOT LOADING THE SCREWS. The 1in. Mirror Mount was NOT in contact with the Base Plate. THE SLOP IN THE CLEARANCE BORES DID NOT PERMIT PRECISE CENTERING AT THIS STAGE.
  2. The set screws were loosened and retracted into the threaded bores to prevent any possible contact with the barrel of the mirror.
  3. The 1in. Mirror Mount was bolted to the Base plate using the M2.5 x 0.45 SHCS with tightening conducted in stages, so that each tab was loaded approximately 1 mm at a time. A final torque of 5 in*lb was applied, per T1100066 (interpolation between #2-56 and #4-40).

Note that the mirror used in this experiment was a flat mirror removed from PMC 10. See item 36 of D1001955-v2 Assembly Drawing.

Evaluation

Using the Fit Test Interface Assy D1600398 and a 36 ounce Jonard Compression/Tension Force Guage, the holding force of the 906919-02 Mirror Mount Spring Plate was tested.

This holding force was found to hold the optic tightly without macroscopic translation in the nominal mount configuration (0.050” axial deflection on spring plate) until acted upon by approximately 6 ounces of force (converted, about 1.7 newtons). This force was applied approximately perpendicular to points on the barrel of the mirror near the face opposite the 3-point contact plane.

SEE 50thou Holding Force.MOV

In contrast to the nominal 0.050” of deflection on the spring plate, under an axial deflection of 0.030” approximately 4 ounces of force (1.1 newtons) were required to overcome the static friction and induce macroscopic translation. This set-up utilized precise shims to reduce the deflection, as shown below.

SEE 30thou Holding Force.MOV

The force required to induce macroscopic translation without the spring plate installed was too small to be measured with the same set-up.

SEE NoSpring Holding Force.MOV

The weight of the optic contributes a force of 1.6E-2 newtons in the nominal orientation, using a calculated volume of 0.78 cm^3 and an estimate density of 2.2 grams per cm^3. The optic was set in the Interface Assy for 2 hours under the nominal configuration, and no gravitational sag was observed (using a pin gage as a probe).

 

Evaluation of Shipping Configuration

In order to push the limits of the holding force of the mount for shipping purposes, the mount was carefully subjected to a repeated jarring motion (held with one hand and "clapped" firmly into the other) with the set screws driven tightly into the barrels. The contact provided by the set screws and the spring plate in conjunction was sufficient that the optic did not macroscopically move (using a pin gage as a probe). Meanwhile, when the optic was held only by the spring plate and the mount was subjected to the same jarring motion, the optic did macroscopically move. The set screw contact on the barrel is very much necessary for shipping of the optic in place within this mount in order to provide sufficient static friction to rigidly couple the optic to the mount.

 

 

 

Attachment 1: image3.JPG
image3.JPG
Attachment 2: image2.JPG
image2.JPG
Attachment 3: image1.JPG
image1.JPG
Attachment 4: image1.JPG
image1.JPG
Attachment 5: image5.JPG
image5.JPG
Attachment 6: image11.JPG
image11.JPG
Attachment 7: 30thou_Holding_Force.MOV
Attachment 8: 50thou_Holding_Force.MOV
Attachment 9: NoSpring_Holding_Force.MOV
Attachment 10: image21.JPG
image21.JPG
Attachment 12: image22.JPG
image22.JPG
  20   Thu Oct 13 14:36:14 2016 AlenaProgressThomasSR3 actuator test

Goal: Position dependent measurement of the temperature at the plate.

Setup: thermocouple next to the center with capton tape

Current: 190, 250 and 300 mA

Expected max temperature: 200 C

Pressure: 5 10-6 torr

I moved the thermocouple from the center because it's washer changes the amount of compression and the measurement it difficult to compare with other once. Also the bump at the washer makes holes in the plating if everything is tighten. The measurement shows that more compression is needed (less washers)

  19   Thu Oct 13 13:12:50 2016 Alena, Rich and LuisElectronicsPZT jitter experimentPZT transfer function

Mesured transfer function of the PZT with a small 1" mirror

Attachment 2: Pitch_TF_Phase.png
Pitch_TF_Phase.png
  18   Tue Oct 11 13:35:00 2016 AlenaProgressThomasSR3 actuator test

Goal: comparison of the temperature at the center of the gold plated plate during two runs: with 4 and with 3 washers (part #4 per D1500387)

Setup: thermocouple at the center, In gasket

Current: 190(207), 250(258) and 300(308) mA

Expected max temperature: 200 C

Pressure: 4 10-7 torr

Using 3 washers instead of 4 improves the thermal contact. The temperature at the thermocouple did not change but the ring heater doesn't warm up as hot as before.

Attachment 2: less_washers_test.png
less_washers_test.png
  17   Thu Oct 6 10:56:01 2016 AlenaProgressThomasSR3 actuator test

Goal: comparison of the temperature at the center of the gold plated plate during two runs: without gasket and with In gasket

Setup: thermocouple at the center

Current: 190, 250 and 300 mA

Expected max temperature: 200 C

Pressure: 2 10-6 torr

In pad under the thermocouple washer after the first run :

Mounting the indium gasket

Results: This two runs did not demonstrate the same dramatic improvement by adding the gasket as previously. The most probable reason is the decreased compression because of adding a thermocouple with a washer to the assembly. Another confirmation of the bad compression is higher ring heater temperature comparing to previous runs.

 

Attachment 1: unnamed.jpg
unnamed.jpg
Attachment 6: repetability_test.png
repetability_test.png
  16   Mon Oct 3 17:06:10 2016 AlenaElectronicsPZT jitter experimentWiring of the QPD

We bought a thorlads QPD to measure the transfer function of the PZT actuator. today I wired a box to connect the QPD to a power supply and to read pich, yaw and the sum.

Attachment 1: pin_diagram.png
pin_diagram.png
  15   Mon Oct 3 07:59:46 2016 AlenaProgressThomasSR3 actuator test

Goal: cheking contact at thermo couple is ok ore needs improovement

Setup: no gasket at the ring heater; indium thermal pad on the thermocouple, thermocoupel at the center of the back plate

Current: 189(207), 250(258), 300 mA - 7:30 AM, 11:30 AM, 3:30 PM

Expected max temperature: 220 C

Pressure: 2 10-7 torr

Results: The temperature measured with the thermocouple almost did  not change by adding indium foil between the washer and the gold plated plate (red curve). Hovewer the temperature at the ring heater changed. A second run with no indium (after reassembling the setup) shown a consistent temperature at the gold plated surface and again completely different temperature at the ring heater. Pressure was the same during all three measurement.

  14   Fri Sep 30 12:16:23 2016 AlenaProgressLaser DamageViewport cover

The modified Viewport cover was installed using the new adapter ring. No leaks. Pump down looked Ok.

 

Attachment 3: 2.jpg
2.jpg
ELOG V3.1.3-